# Standard Gaussian Process is All You Need for High-Dimensional Bayesian Optimization

### v2

@@ 3.2 Structural Assumption in Input Space
Low-dimensional Embedding. Wang et al. (2016) proposed a method named REMBO, building the GP surrogate in a low-dimensional embedding space, $\mathcal{Z}\subset\mathbb{R}^{d_{\text{emb}}}$ where $d_{\text{emb}}\ll d$. The acquisition function is optimized in the embedding space. When querying the function value, the input is recovered by ${\bf x}={\bf A}{\bf z}$ where ${\bf z}\in\mathcal{Z}$, and ${\bf A}$ is a random matrix. If the recovered ${\bf x}$ is out of the domain $\mathcal{X}$, it will be clipped as a boundary point. To avoid clipping, Nayebi et al. (2019) proposed HESBO, which randomly duplicates ${\bf z}$ (and/or flip the signs) to form ${\bf x}$. This way, the optimization of ${\bf z}$ can respect the same bounding constraints as in $\mathcal{X}$. This essentially impose further constraints on the embedding space. A more flexible solution, ALEBO, was proposed in  (Letham et al., 2020). ALEBO also uses a random matrix ${\bf A}$ to recover the input ${\bf x}$. But when maximizing the acquisition function, ALEBO incorporates a constraint that ${\bf A}{\bf z}$ must be in $\mathcal{X}$, thereby avoiding the clipping issue.
Partial Variable Dependency. Another strategy is to trim down the input variables. The recent start-of-the-art, SaasBO (Eriksson and Jankowiak, 2021), uses a horse-shoe prior (Carvalho et al., 2009) over the length-scale parameter for each variable. As the horse-shoe prior induces strong sparsity, a massive number of variables can be pruned, substantially reducing the input dimension. For effective training with the horse-shoe prior, SaasBO used a No-U-Turn Sampler (NUTS) (Hoffman et al., 2014). While one can also use MAP estimation, the performance is often much inferior.
@@ 4 Comprehensive Evaluation
We now examine and analyze the performance of standard BO in high-dimensional optimization. To this end, we conducted a series of systematic investigations.
@@ 4.1 Experimental Settings
We employed twelve benchmarks, of which six are synthetic and the other six are real-world problems. For clarity in the discussion, all the tasks aim for function maximization.
Synthetic Benchmarks. We considered four popular synthetic functions: Ackley, Rosenbrock, Hartmann6, and Stybtang. Definitions of these functions are provided in Section A.1 of the Appendix. Each task is represented by $\text{Fun}(d,d^{\prime})$, where $d$ is the input dimension that the BO methods optimize for, and $d^{\prime}$ is the number of effective variables ($d^{\prime}\leq d$). The target function is computed using the first $d^{\prime}$ variables. The benchmarks and the structures within the target functions are summarized in Table 1.
@@
  • Mopta08 (124)  (Jones, 2008), a vehicle design problem. The objective is to minimize the mass of a vehicle with respect to 124 design variables, subject to 68 performance constraints. We followed (Eriksson and Jankowiak, 2021) to encode the constraints as a soft penalty, which is added into the objective.
  • SVM (388) (Eriksson and Jankowiak, 2021), a hyper-parameter tuning problem that optimizes 3 regularization parameters and 385 length-scale parameters for a kernel support vector machine regression model.
  • Rover (60), a rover trajectory planning problem from (Wang et al., 2018). The goal is to find an optimal trajectory determined by the location of 30 waypoints in a 2D environment.
  • DNA (180) (vSehić et al., 2022), a hyper-parameter optimization problem that optimizes 180 regularization parameters for weighted lasso regression on an DNA dataset (Mills, 2020). Prior analysis (vSehić et al., 2022) shows that only around 43 variables are relevant to the target function.
  • NAS201 (30)  (Dong and Yang, 2020), a neural architecture search problem on CIFAR-100 dataset.
  • NN-Prune (151). We provided an additional real-world benchmark on neural network pruning (Blalock et al., 2020). We considered ResNet-152 (He et al., 2015), a deep residual network with 151 convolution layers and one linear layer, including more than 60 million parameters in total. We introduced a pruning rate $x_{i}$ for each convolution layer $i$. Given a global pruning ratio set to 0.5, the objective is to optimize the per-layer pruning rates to maximize the prediction accuracy on a test dataset after pruning and fine tuning. The details are given in Section A.2 of the Appendix.
@@
Table 2: BO methods and their structural assumption.
Methods. To execute the standard BO, we used GP regression with Matérn-5/2 ARD kernel (see (3)). We also tested with Matérn-5/2 kernel with a single length-scale parameter. We denoted these two choices as GP-ARD and GP, respectively. Throughout the experiments, we used UCB as the acquisition function where the exploration level $\lambda$ was set to 1.5. We used GPyTorch (Gardner et al., 2021) for surrogate training and BOTorch (Balandat et al., 2020) for acquisition function computation.
@@
We compared with six state-of-the-art high-dimensional BO methods, including Tree UCB (Han et al., 2020), RDUCB (Ziomek and Bou-Ammar, 2023), HESBO (Nayebi et al., 2019), ALEBO (Letham et al., 2020), SaasBO (Eriksson and Jankowiak, 2021), and TURBO (Eriksson et al., 2020). The first five have been introduced in Section 3. TURBO is designed for a special scenario where the target function can be extensively evaluated (e.g., tens of thousands evaluations). This is not common in BO applications, as the function evaluation is typically deemed expensive, and one aims to evaluate the target function as few times as possible. TURBO searches for trust-regions in the input domain, learns a set of local GP models over each region, and then uses an implicit bandit strategy to decide which local optimization runs to continue. Table 2 summarizes the structural assumptions made by all the methods. We used the original implementation and default (recommended) settings for the competing methods. See details in Appendix Section A.3.
@@
For each optimization task, we randomly queried the target function for 20 times to generate the initial data. We then ran each method for 400 steps for synthetic tasks and for 300 steps for real-world tasks. We tested TURBO with a single trust region and five trust regions, denoted by TURBO-1 and TURBO-5 respectively. For HESBO and ALEBO, we varied $d_{\text{emb}}\in\{10,20\}$ and denote the choice as HESBO/ALEBO-{10, 20}. For SaasBO, we performed tests using both NUTS and MAP estimation for surrogate training, denoted by SaasBO-NUTS and SaasBO-MAP, respectively. For standard BO, we trained the GP regression by maximizing the marginal likelihood using ADAM optimization (Kingma and Ba, 2014) (learning rate $0.1$ and 500 steps, as suggested in GPyTorch documents). We treated the mean function as a constant and jointly estimated it with the kernel parameters and noise variance.
4.2 Results and Discussion
4.2.1 Optimization Performance
@@
We present the ultimate optimization outcomes for synthetic benchmarks in Fig. 1 and for real-world benchmarks in Fig. 2. These are violin plots that illustrate the average of the obtained maximum function value over ten runs (the middle bar), the best (the top bar), the worst (the bottom bar), and the distribution of the obtained maximum (the shaded region). We can observe that on average, the standard BO with GP-ARD outperforms the competing high-dimensional BO methods in optimization every target function, often by a large margin, except that in the cases of Hartman6(300, 6) and NN-Prune, the performance with GP-ARD is slight worse (but still very close to) the best baseline(s).
@@
It is noteworthy that the performance of the competing methods appears to be sensitive to the structure within the target function. For example, SaasBO imposes a strong inductive bias that only a small number of variables is related to the objective (via the usage of the horse-shoe prior). The target function in Hartman6(300, 6) and DNA aligns well with this assumption. As a result, SaasBO exhibits outstanding performance in both tasks. However, when the objective deviates from this assumption, as seen in cases like Ackley(150, 150), StybTang(200, 200), NN-Prune, and even Rosenbrock(300, 100), the performance of SaasBO experiences a dramatic drop. Another example is RDUCB and Tree-UCB, which assume that the target function can be decomposed as a summation of a group of functions, each over a small number of variables. RDUCB and Tree-UCB perform well in synthetic benchmarks that match this assumption, such as Rosenbrock(100, 100), Rosenbrock(300, 100), and StybTang(200, 200). However, their performance drops significantly in cases that do not align with their assumption, such as Ackley(150, 150) and Ackley(300, 150).
@@
In contrast, GP-ARD exhibits much greater robustness to the structure of the target function. The standard BO with GP-ARD consistently ranks among the best in all tasks, regardless of whether the target function depends only on a subset of variables, can be decomposed, etc. Interestingly, the standard BO shows superior performance even when using a single length-scale parameter for GP, which closely matches GP-ARD across most tasks, except for Ackley(300, 150), GP struggled in finding right locations to query, leading to relative poor performance.
@@
We also examined the runtime performance of each method. We report the maximum function value obtained throughout the optimization process in Fig. 5 and Fig. 6 of the Appendix. It is noticeable that after a few initial steps, the standard BO with GP-ARD (and also with GP in most cases) consistently generated superior queries, finding larger function values. Its curves are mostly above those of the competing methods, indicating a more efficient optimization of the target function. This trend is particularly evident in examples such as Rosenbrock(100, 100), StybTang(200, 200), Mopta08, Rover, SVM, and more.
@@
The collective results demonstrate that the standard GP with ARD kernels is already effective for high-dimensional optimization. In comparison with using more complex surrogate models with additional assumptions, GP-ARD exhibits greater flexibility in adapting to the target function whose intrinsic structures are often agnostic in practice, and provides superior optimization results.
@@
4.2.2 Surrogate Learning
Next, we explicitly evaluated the performance of the standard GP in learning high-dimensional functions to confirm if it can serve as a reliable surrogate model for high-dimensional BO. We sampled 300 training examples from all the six synthetic benchmark functions and four real-world benchmark tasks, including Mopta08 (124), Rover (60), DNA (180) and SVM (388). We sampled another 100 examples for testing. We compared with the surrogate models used in SaasBO and ALEBO (with $d_{\text{emb}}=10$ and $d_{\text{emb}}=20$). SaasBO was trained by NUTS. We repeated the experiment for five times, and evaluated the average test log-likelihood (LL), mean-square-error (MSE), and the standard deviation.
The results are reported in Table 3 and Appendix Table 5. We can see that in the vast majority of cases, GP-ARD achieves the largest test LL and the smallest MSE. GP-ARD and GP together nearly always outperforms SaasBO, ALEBO-10 and ALEBO-20. Notably, GP-ARD surpasses ALEBO by a substantial margin, while SaasBO exhibits a closer prediction accuracy. In addition, we examined the alignment between the prediction and the ground-truth across the test examples from Mopta and DNA. As illustrated in Fig. 4, the prediction (as well the predictive standard deviation) of GP-ARD matches the ground-truth equally as well as SaasBO-NUTS, with a slight improvement over GP (using a single length-scale parameter).
Together these results demonstrate that directly incorporating high-dimensional inputs into the kernel function does not necessarily compromise the learning performance. In other words, this implies that, even when the input space is high dimensional, commonly used kernels (e.g., the Matérn-5/2) remains expressive enough to capture the similarities in the space, allowing them to perform well in covariance modeling. On the contrary, enforcing (unnecessary) low-rank structures or sparsity might cause information loss, potentially detrimental to the learning performance. Our observation is consistent with the recent work in kernel operator learning (Batlle et al., 2024), where the standard kernel (GP) regression has demonstrated the state-of-the-art performance in learning function-to-function mappings. In this context, the input comprises of a discretized function, typically of very high-dimensionality (e.g., several thousands).
4.2.3 Computational Efficiency
So far, in our BO loop, the GP model is trained via maximum likelihood estimation (MLE). Although the performance has been promising, it is worth investigating whether a full Bayesian approach would yield substantial improvements. In particular, we can use MCMC to perform posterior inference of the GP parameters, and then integrate them out during the prediction. We conducted another round of BO experiments. For each step, we used NUTS to train GP-ARD. We adopted the same settings as used in SaasBO-NUTS (Eriksson and Jankowiak, 2021), where the mean parameter is assigned the standard Gaussian prior $\mathcal{N}(0,1)$, amplitude $\text{Gamma}(2,0.15)$, and noise variance $\text{Gamma}(0.9,10)$. We then assigned a uniform prior over each length-scale, $\text{Uniform}(0.001,30)$. The number of warm up steps was set to 512, and after the warm up stage, we ran 256 steps, and selected the samples at every 16-th step as the posterior samples. We show the running and final optimization performance in Fig 3 , and Appendix Fig. 7. It can be observed that while NUTS contributes to a slight improvement in several tasks, the enhancement is very marginal. However, running NUTS is much more computationally costly compared to MLE. For example, on NAS201 and DNA, NUTS takes 68x and 40x running time than MLE in each optimization step. Therefore, for practical purposes, using maximum likelihood estimation (MLE) for Bayesian Optimization with GP-ARD appears sufficient. It is computationally efficient, and the sacrifice of optimization performance is minimal compared to much more costly training methods like NUTS.
By contrast, for more complex surrogate models, such as SaasBO, the MCMC based training might be more crucial for achieving good performance. As depicted in Fig. 2, SaasBO-NUTS demonstrates a substantial improvement over SaasBO-MAP across all real-world benchmarks, excluding NAS201. However, the computational cost of running SaasBO-NUTS is considerably higher, ranging from 5.4x (NAS201) to 33.9x (SVM, DNA) compared to the running time of SaasBO-MAP. This substantial increase in computational requirements may pose a severe bottleneck for practical usage. Detailed per-optimization-step running time for these methods are provided in Table 4 and Appendix Table 6.
Task GP-ARD GP SaasBO ALEBO-10 ALEBO-20
Mopta08(124) 0.613$\pm$0.156 -0.198$\pm$0.233 0.584$\pm$0.158 -1.176$\pm$0.080 -1.177$\pm$0.327
Rover(60) -1.209$\pm$0.194 -0.908$\pm$0.068 -0.909$\pm$0.108 -1.270$\pm$0.094 -1.687$\pm$0.251
DNA(180) 0.005$\pm$0.175 -0.459$\pm$0.142 -0.066$\pm$0.225 -1.346$\pm$0.108 -1.572$\pm$0.432
SVM(388) 0.296$\pm$0.097 -0.090$\pm$0.065 0.160$\pm$0.113 -1.175$\pm$0.116 -1.077$\pm$0.148
Ackley(150, 150) 0.720$\pm$0.092 0.899$\pm$0.058 0.809$\pm$0.053 -0.852$\pm$0.091 -0.266$\pm$0.091
Ackley150(300, 150) 0.617$\pm$0.191 0.367$\pm$0.081 0.536$\pm$0.095 -0.870$\pm$0.097 -0.478$\pm$0.204
Hartmann6 (300, 6) 1.681$\pm$0.110 -1.031$\pm$0.117 2.152$\pm$0.971 -1.433$\pm$0.249 -1.689$\pm$0.411
Rosenbrock(100, 100) 0.602$\pm$0.069 0.172$\pm$0.099 0.553$\pm$0.084 -1.007$\pm$0.121 -0.854$\pm$0.355
Rosenbrock(300, 100) 0.436$\pm$0.098 -0.256$\pm$0.071 0.265$\pm$0.115 -1.118$\pm$0.065 -1.380$\pm$0.329
Stybtang(200, 200) 0.558$\pm$0.122 0.215$\pm$0.076 0.463$\pm$0.098 -1.050$\pm$0.099 -1.238$\pm$0.436
Table 3: Test log likelihood of each surrogate model. The results were averaged from five runs.
Mopta08 (124) Rover (60) NAS201 (30) DNA (180) SVM (388)
GP-ARD 11.23 9.84 5.76 14.34 32.16
GP-ARD-NUTS 492.62 404.17 393.62 575.03 768.29
SaasBO-MAP 46.53 41.36 42.35 51.33 67.37
SaasBO-NUTS 985.76 375.06 229.07 1739.56 2269.50
Figure 3: Optimization performance when using GP-ARD trained with maximum likelihood estimation (MLE) and NUTS on SVM (388) and DNA (180). The results were summarized from 10 runs.
(a) GP-ARD
(b) GP

### v3

@@ 3.2 Structural Assumption in Input Space
Low-dimensional Embedding. Wang et al., [32] proposed a method named REMBO, building the GP surrogate in a low-dimensional embedding space, $\mathcal{Z}\subset\mathbb{R}^{d_{\text{emb}}}$ where $d_{\text{emb}}\ll d$. The acquisition function is optimized in the embedding space. When querying the function value, the input is recovered by ${\bf x}={\bf A}{\bf z}$ where ${\bf z}\in\mathcal{Z}$, and ${\bf A}$ is a random matrix. If the recovered ${\bf x}$ is out of the domain $\mathcal{X}$, it will be clipped as a boundary point. To avoid clipping, Nayebi et al., [25] proposed HESBO, which randomly duplicates ${\bf z}$ (and/or flip the signs) to form ${\bf x}$. This way, the optimization of ${\bf z}$ can respect the same bounding constraints as in $\mathcal{X}$. This essentially imposes further constraints on the embedding space. A more flexible solution, ALEBO, was proposed in  [18]. ALEBO also uses a random matrix ${\bf A}$ to recover the input ${\bf x}$. But when maximizing the acquisition function, ALEBO incorporates a constraint that ${\bf A}{\bf z}$ must be in $\mathcal{X}$, thereby avoiding the clipping issue.
Partial Variable Dependency. Another strategy is to trim down the input variables. The recent start-of-the-art, SaasBO [6], uses a horse-shoe prior [4] over the length-scale parameter for each variable. As the horse-shoe prior induces strong sparsity, a massive number of variables can be pruned, substantially reducing the input dimension. For effective training with the horse-shoe prior, SaasBO used the No-U-Turn Sampler (NUTS) [11]. While one can also use MAP estimation, the performance is often much worse.
@@ 4 Comprehensive Evaluation
We now examine and analyze the performance of standard BO in high-dimensional optimization. To this end, we conducted a series of systematic experiments.
@@ 4.1 Experimental Settings
We employed eleven benchmarks, of which six are synthetic and five are real-world problems. For clarity in the discussion, all the tasks aim for function maximization.
@@
Synthetic Benchmarks. We considered four popular synthetic functions: Ackley, Rosenbrock, Hartmann6, and Stybtang. Definitions of these functions are provided in Section A.1 of the Appendix. Each task is represented by “$\text{Fun}(d,d^{\prime})$”, where $d$ is the input dimension that BO methods optimize for, and $d^{\prime}$ is the number of effective variables ($d^{\prime}\leq d$). The ground-truth of the target function values is computed with the first $d^{\prime}$ variables. The benchmarks and the structures within the target functions are summarized in Table 2.
@@
  • Mopta08 (124)  [13], a vehicle design problem. The objective is to minimize the mass of a vehicle with respect to 124 design variables, subject to 68 performance constraints. We followed [6] to encode the constraints as a soft penalty, which is added into the objective.
  • SVM (388) [6], a hyper-parameter tuning problem that optimizes 3 regularization parameters and 385 length-scale parameters for a kernel support vector machine regression model.
  • Rover (60), a rover trajectory planning problem from [31]. The goal is to find an optimal trajectory determined by the location of 30 waypoints in a 2D environment.
  • DNA (180) [35], a hyper-parameter optimization problem that optimizes 180 regularization parameters for weighted lasso regression on an DNA dataset [22]. Prior analysis [35] shows that only around 43 variables are relevant to the target function.
  • NAS201 (30)  [5], a neural architecture search problem on CIFAR-100 dataset.
@@
Table 3: BO methods and their structural assumption about the target function.
Methods. To execute the standard BO (SBO), we used GP regression with ARD Matérn-5/2 kernel (see (1)). We used UCB as the acquisition function where the exploration level $\lambda$ was set to 1.5. We used GPyTorch [9] for surrogate training. To compare with other high-dimensional BO methods, we considered two training schemes. One performs MAP estimation with LBFGS. The maximum number of iterations was 15K and tolerance level 2E-9. We employed a diffuse prior over each length-scale, i.e., Uniform(0.001, 30). The noise variance was assigned a diffuse Gamma prior, Gamma(1.1, 0.05), and the amplitude the prior Gamma(2.0, 0.15). We denote this method as SBO-1. The second training scheme simply uses ADAM optimization [17] to perform MLE, with learning rate $0.1$ and the maximum number of epochs was set to $500$. We denote this method as SBO-2. We also tested with Matérn-5/2 kernel with a single length-scale parameter, which we denoted as SBO-3. We used BOTorch [2] for acquisition function maximization, which uses LBFGS. Additional evaluations on alternative choices of the kernel, training scheme, and acquisition function are provided in Appendix Section B.
@@
We compared with six state-of-the-art high-dimensional BO methods, including Tree UCB [10], RDUCB [34], HESBO [25], ALEBO [18], SaasBO [6], and TURBO [7]. The first five have been introduced in Section 3. TURBO is designed for a special scenario where the target function can be extensively evaluated, e.g., tens of thousands evaluations. This is uncommon in BO applications, as the function evaluation is typically deemed expensive, and one aims to evaluate the target function as few as possible. TURBO searches for trust-regions in the input domain, learns a set of local GP models over each region, and then uses an implicit bandit strategy to decide which local optimization runs to continue. We used the original implementation and default (recommended) settings for the competing methods. In addition, we compared with two BO methods based on Bayesian neural networks (BNN). One uses Laplace approximation and the other Hamiltonian Monte-Carlo (HMC) sampling for posterior inference. We used a high-quality open-source implementation for each BNN based method. Table 3 summarizes the structural assumptions made by all the methods. The details of all the competing methods are given in Appendix Section A.2.
@@
For each optimization task, we randomly queried the target function for 20 times to generate the initial data. We then ran each method for 400 steps for synthetic tasks and for 300 steps for real-world tasks. We tested TURBO with a single trust region and five trust regions, denoted by TURBO-1 and TURBO-5 respectively. For HESBO and ALEBO, we varied $d_{\text{emb}}\in\{10,20\}$ and denote the choice as HESBO/ALEBO-{10, 20}. For SaasBO, we performed tests using both NUTS and MAP estimation for surrogate training, denoted by SaasBO-NUTS and SaasBO-MAP, respectively. For a reliable comparison, we ran every method for ten times.
4.2 Optimization Performance
@@
Ultimate Outcomes. We first examined the ultimate optimization outcomes of each method. The results are reported in Fig. 1 for synthetic benchmarks and in Fig. 2 for real-world benchmarks. Following [6], we used violin plots that illustrate the average of the obtained maximum function value over the ten runs (the middle bar), the best (the top bar), the worst (the bottom bar), and the distribution of the obtained maximum (the shaded region). We can observe that on average, the standard BO, including both SBO-1 and SBO-2, outperforms the competing methods at every target function, often by a large margin, except that in the cases of Hartman6(300, 6) and NAS201, the performance of SBO is slight worse than (but still very close to) the best baseline(s).
@@
It is noteworthy that the performance of the competing high-dimensional BO methods appears to be sensitive to the structure within the target function. For example, SaasBO imposes a strong inductive bias that only a small number of variables is related to the objective (via the usage of the horse-shoe prior). The target function in Hartman6(300, 6) and DNA aligns well with this assumption. As a result, SaasBO exhibits outstanding performance in both tasks. However, when the objective deviates from this assumption, as seen in cases like Ackley(150, 150), StybTang(200, 200) and even Rosenbrock(300, 100), the performance of SaasBO experiences a dramatic drop. Another example is RDUCB and Tree-UCB, which assume that the target function can be decomposed as a summation of a group of functions, each over a small number of variables. RDUCB and Tree-UCB perform well in synthetic benchmarks that match this assumption, such as Rosenbrock(100, 100), Rosenbrock(300, 100), and StybTang(200, 200). However, their performance drops significantly in cases that do not align with their assumption, such as Ackley(150, 150) and Ackley(300, 150). Across all the benchmarks, the BNN based methods performed worse than most GP based methods. This might be due to that under low data regimes, it is more challenging for a neural network to capture the shape of the target function and quantify the uncertainty, especially for high-dimensional targets.
@@
In contrast, the standard BO exhibits much greater robustness to the structure of the target function. Using standard GP equipped with Matérn kernels, both SBO-1 and SBO-2 consistently rank among the best in all tasks, regardless of whether the target function depends only on a subset of variables, can be decomposed, etc. Interestingly, as shown in Appendix Fig.7, even using a single length-scale parameter, the standard BO, i.e., SBO-3, still consistently shows superior performance, which closely matches ARD Matérn across most tasks, except for Ackley(300, 150), SBO-3 struggled in finding right locations to query, leading to poor performance.
Runtime Performance. Next, we examined the runtime performance. For each method, we report the average maximum function value obtained at each step in Fig. 5 and Fig. 6 of the Appendix. We supplement the runtime standard deviation and the result of SBO-3 in Fig. 8 of the Appendix. Note that since we have twelve baseline methods, reporting the standard deviations of all the methods makes the figure very cluttered. To ensure clarity, in Fig. 8, we show the standard deviation of the best, worst, and median performed baselines so far at each step.
@@
One can see that after a few initial steps, SBO-1 and SBO-2 (and SBO-3 in most cases) consistently generated superior queries, finding larger function values. Their curves are mostly above those of the competing methods, indicating a more efficient optimization of the target function. This trend is particularly evident in examples such as Rosenbrock(100, 100), StybTang(200, 200), Mopta08, Rover, SVM, and more.
@@
The collective results demonstrate that the standard GP (with Matérn kernels) is effective for high-dimensional optimization. In comparison with more complex surrogate models using additional assumptions, the standard GP exhibits greater flexibility in adapting to the target function whose intrinsic structures are often agnostic in practice, and enables superior optimization results.
@@
4.3 Surrogate Learning
Next, we explicitly evaluated the performance of the standard GP in learning high-dimensional functions to confirm if it can serve as a reliable surrogate model for high-dimensional BO. To this end, from the optimization trajectories of all the methods, we randomly sampled 300 training examples from the six synthetic benchmark functions and four real-world benchmark tasks, including Mopta08 (124), Rover (60), DNA (180) and SVM (388). We sampled another 100 examples for testing. We compared with the surrogate models used in SaasBO, namely, placing a horse-shoe prior over the the length-scale parameters to prune the variables. We also compared with a random projection GP, denoted as ProjGP, to align with the baselines employing low-dimensional embeddings. ProjGP employs the kernel function in the form of $\vartheta({\bf x},{\bf x}^{\prime})=\kappa({\bf A}{\bf x},{\bf A}{\bf x}^{\prime})$, where ${\bf A}$ is a random projection matrix. We varied the rank of ${\bf A}$ from 10 and 20. Each element is sampled from the standard normal distribution. In addition, we compared with an additive GP, denoted as AddGP, where the kernel takes the form $\vartheta({\bf x},{\bf x}^{\prime})=\sum_{j}\kappa({\bf x}^{j},{\bf x}^{\prime j})$ where each ${\bf x}^{j}\subset{\bf x}$ and ${\bf x}^{\prime j}\subset{\bf x}^{\prime}$, ${\bf x}$ and ${\bf x}^{\prime}$ are the union of ${\bf x}^{j}$’s and ${\bf x}^{\prime j}$’s, respectively. AddGP imposes a decomposition structure in the functional space; see (5). We tested AddGP with randomly partitioning the input into 10 variables per group. We tested two versions: one using non-overlapping partitions, and the other overlapping partitions with one variable shared by successive groups. We trained the standard GP with ARD Matérn-5/2 kernel and with a single length-scale Matérn-5/2 kernel using MLE. SaasBO was trained by NUTS. For a fair comparison, ProjGP and AddGP were based on ARD Matérn-5/2 and trained via MLE as well.
(a) ARD Matérn
(b) Matérn